If a product sells for $22, variable costs are $12 and fixed costs are $300,000, what would total sales have to be in order to break-even?

Multiple Choice

  • $3,000,000

  • $660,000

  • $300,000

  • $360,000
Answer #1

Contribution margin=Sales-Variable cost

=22-12=$10 per unit

At breakeven;Total contribution margin=Total fixed costs=$300,000

Hence sales units=(300,000/10)

=30,000 units.

=(30,000 units*$22)

=$660,000